OpenStudy (colcaps):

The sum of the interior angles of a convex polygon is 3060°. Find the number of sides and identify the type of the polygon

OpenStudy (colcaps):

last question

OpenStudy (colcaps):

so 3060/180?

OpenStudy (anonymous):

(n-2)180=3060 n is the number of sides

ganeshie8 (ganeshie8):

that gives u (n-2)

OpenStudy (colcaps):

but the sides arent given yet

ganeshie8 (ganeshie8):

u familiar wid formula given by @AbhimanyuPudi right

OpenStudy (colcaps):

yeah thats the formula i used earlier

ganeshie8 (ganeshie8):

yeah, in the formula, (n-2)*180 = sum of interior angles

ganeshie8 (ganeshie8):

n is the # of sides

ganeshie8 (ganeshie8):

which we need to find out

OpenStudy (colcaps):

so ill divide 3060/180?

ganeshie8 (ganeshie8):

thats right ! (n-2)*180 = 3060 n-2 = 3060/180

OpenStudy (colcaps):

so (17-2)*180=2700

OpenStudy (colcaps):

oh wait

ganeshie8 (ganeshie8):

NO

OpenStudy (colcaps):

sorry i forgot to divide

ganeshie8 (ganeshie8):

ok.. :)

ganeshie8 (ganeshie8):

@waterineyes :)

OpenStudy (colcaps):

15 17 15 17

OpenStudy (anonymous):

What happened @ganeshie8 ??

OpenStudy (colcaps):

15=17

OpenStudy (anonymous):

When I last logged in at 2.50 pm , you were at 92 and now 93.. So quick.. Anyways, congratulations for Green..

ganeshie8 (ganeshie8):

ohh i was about to greet u, then u logged out, looks u too busy wid college :)

OpenStudy (anonymous):

@Colcaps you will get n-2 = 3060/180 = 17 So, n-2=17 then n=?

OpenStudy (colcaps):

15

OpenStudy (anonymous):

Yes that's the answer..the number of sides

OpenStudy (anonymous):

\(\color{green}{\large \frak{Congratulations \; \; For \; \; green \; \;@ \; \; ganesh..}}\)

OpenStudy (colcaps):

thansk @AbhimanyuPudi and @ganeshie8

OpenStudy (anonymous):

congratulations from my side too @ganeshie8

OpenStudy (anonymous):

welcome:) @Colcaps

ganeshie8 (ganeshie8):

thanks @waterineyes :)

ganeshie8 (ganeshie8):

hey check ur arithmetic again...

ganeshie8 (ganeshie8):

@Colcaps

OpenStudy (anonymous):

Sorry dude..its n=19

OpenStudy (anonymous):

Yes I am busy now a days in college.. I don't have sufficient time to be here on this site... I will be back but all I can do is wait for the free time..

OpenStudy (colcaps):

19?

OpenStudy (colcaps):

wait did the sign transpose?

OpenStudy (anonymous):

Yes, n-2=17 Add 2 on both sides n=17+2=19

ganeshie8 (ganeshie8):

yeah il msg u i guess :)

OpenStudy (colcaps):

so the polygon is called enneadecagon?

OpenStudy (anonymous):

A Polygon with 19 sides is called ...... ?? I don't know...

OpenStudy (anonymous):

yes you are right...you can also say nonadecagon

OpenStudy (anonymous):

Yeah it is what you said..